OpenStudy (anonymous):

How do you convert y=2(x-5)(x+2) into standard form, y=ax^2+bx+c?

OpenStudy (anonymous):

Multiply it.

OpenStudy (anonymous):

so would i do 2(x-5) 2(x+2) ?

OpenStudy (anonymous):

hi

OpenStudy (gorv):

2(x^2-5x+2x-10)=2x^2-2*3x-10=0 2x^2-6x-10

OpenStudy (gorv):

y=2x^2-6x-10 @ali1029

OpenStudy (anonymous):

y=2(x-5)(x+2) y=(2x-10)(x+2) y=2x(x+2)-10(x+2) \[y=2x^2+4x-10x-20 =2x^2-6x-20\] which is in the form of y=ax^2+bx+c @ali1029

OpenStudy (anonymous):

so it would be 2x^2-6x-20 and not 2x^2-6x-10 ?

OpenStudy (anonymous):

Yes, it would be 2x^2-6x-20 @ali1029

OpenStudy (ranga):

2x^2−6x−20

OpenStudy (anonymous):

thank you so much for explaining and helping! :)